OpenStudy (anonymous):

Please help. Integrate sinh^2(x) cosh^2(x) dx

OpenStudy (anonymous):

Help pls.

OpenStudy (dumbcow):

For this one i would rewrite in exponential form ---> http://en.wikipedia.org/wiki/Hyperbolic_function#Standard_algebraic_expressions changing the integral to \[\int\limits \frac{ e^{-8x} -2 e^{-4x} +1}{16 e^{-4x}} dx\] then use substitution u = e^(-4x) du = -4 e^(-4x) dx \[-\frac{1}{64} \int\limits \frac{u^2 -2u +1}{u^2}\] \[ = -\frac{1}{64}( u -2 \ln u - \frac{1}{u})\] plugging back in and simplifying \[-\frac{1}{64} (\frac{e^{-8x} -1}{e^{-4x}} + 8x)\] distribute the neg and put back in hyperbolic form \[= \frac{1}{32}( \sinh (4x) - 4x)\]

OpenStudy (anonymous):

Thank you very much sir

OpenStudy (anonymous):

Another suggestion: integrate by parts, setting \[\begin{matrix} u=\sinh x&&&dv=\sinh x\cosh^2x~dx\\ du=\cosh x~dx&&&v=\frac{1}{3}\cosh^3x \end{matrix}\] So \[\begin{align*}\int\sinh^2x\cosh^2x~dx&=\frac{1}{3}\cosh^3x\sinh x-\frac{1}{3}\int\cosh^4x~dx\\\\ &=\frac{1}{3}\cosh^3x\sinh x-\frac{1}{3}\int\cosh^2x\cosh^2x~dx\\\\ &=\frac{1}{3}\cosh^3x\sinh x-\frac{1}{3}\int(1+\sinh^2x)\cosh^2x~dx\\\\ \frac{4}{3}\int\sinh^2x\cosh^2x~dx&=\frac{1}{3}\cosh^3x\sinh x-\frac{1}{3}\int \cosh^2x~dx\\\\ \int\sinh^2x\cosh^2x~dx&=\frac{1}{4}\cosh^3x\sinh x-\frac{1}{4}\int \frac{\cosh2x+1}{2}~dx\\\\ &=\cdots \end{align*}\]